Which of the following sets are closed in C. (Justify your answer)

a) A={z∈C:Rez<1}∪{z∈C:Imz≥1}

b) B={z∈C:-∞<x≤3}


1
Expert's answer
2021-12-27T18:31:50-0500

a) Let "z_n=1-1/n\in A", since for all "n\in\mathbb{N}" "Re\,z_n<1".

"a=\lim\limits_{n\to\infty}z_n=1\notin A", because "Re\, a=1" and "Im\, a=0<1". Therefore, the set "A" is not closed.

b) Let "z_n" be an arbitrary sequence in "B" converging to a limit "b=\lim\limits_{n\to\infty}z_n".

"|Re\, b-Re\, z_n|=|Re\, (b-z_n)|\leq |b-z_n|\to 0". Then "Re\, b=\lim\limits_{n\to\infty}Re\, z_n\leq 3".

So, "b\in B" and therefore, "B" is a closed set.